# In this paper, we show that the constant-dimensional Weisfeiler–Leman algorithm for groups (Brachter & Schweitzer, LICS 2020) can be fruitfully used to improve parallel complexity upper bounds on isomorphism testing for several families of groups. In particular, we show:

### v6

@@ 7.3 Count-free WL and Abelian groups
Case 3: $\langle g^{\prime}_{1}\rangle\cap A_{1}=1$. In this case, $g^{\prime}_{1}$ is part of a basis for $\langle A_{1},g^{\prime}_{1}\rangle$, thus the type of the latter is either $(a+1,b,c)$ if the order of $g^{\prime}_{1}$ is 2, or $(a,b+1,c)$ if the order of $g^{\prime}_{1}$ is 4. In the former case, we have $a+1\leq n-2$ since $G$ only contains $(\mathbb{Z}/2\mathbb{Z})^{n-2}$ as a direct factor, but not any larger power of $(\mathbb{Z}/2\mathbb{Z})$, that is, $G_{n}$ does not contain subgroups of type $(n-1,*,*)$. (And in any case we must have $a+1<n$ since there are only $n$ pebbles.) Thus $a<n$, and by Lemma 7.5 there is an element $h^{\prime}_{1}\in H$ of order 2 that is not a square in $H$ and is not in $B_{1}$, and Duplicator may respond with $h^{\prime}_{1}$. As before, since $\langle h^{\prime}_{1}\rangle\cap B_{1}=1$, there is a unique extension of $\varphi$ to $\langle A_{1},g^{\prime}_{1}\rangle$ that sends $g^{\prime}_{1}\mapsto h^{\prime}_{1}$ which is also good, and thus the new pebbling is good.
@@
Proposition 7.10 (Order finding in WL-III).
@@
Thm. 7.12 is an example of the strategy of using count-free WL, followed by a limited amount of counting afterwards. (We contrast this with the parallel implementation of the classical (counting) WL algorithm, which—for fixed $k$—uses a polynomial number of Majority gates at each iteration [GV06].) After the fact, we realized this same bound could be achieved by existing techniques; we include both proofs to highlight an example of how WL was used in the discovery process.
@@
Suppose now that $H$ is Abelian. We run the count-free $(O(1),O(\log\log n))$-WL using the parallel WL implementation due to Grohe & Verbitsky. As $G$ and $H$ are non-isomorphic Abelian groups, they have different order multisets. In particular, there exists an order class of greater multiplicity in $G$ than in $H$. By Prop. 7.10, two elements with different orders receive different colors. We use a $\beta_{1}\textsf{MAC}^{0}$ circuit to distinguish $G$ from $H$. Using $O(\log n)$ non-deterministic bits, we guess the color class $C$ where the multiplicity differs. At each iteration, the parallel WL implementation due to Grohe & Verbitsky records indicators as to whether two $k$-tuples receive the same color. As we have already run the count-free WL algorithm, we may in $\textsf{AC}^{0}$ decide whether two $k$-tuples have the same color. For each $k$-tuple of $G^{k}$ having color class $C$, we feed a $1$ to the Majority gate. For each $k$-tuple of $H^{k}$ having color class $C$, we feed a 00 to the Majority gate. The Majority gate outputs a $1$ if and only if there are strictly more $1$’s than 00’s. The result now follows. ∎
@@ 8 Conclusion
It is often the case that if an uncolored class of graphs is identified by WL, then so is the corresponding class of colored graphs. So if constant-dimensional WL identifies a class of graphs, it can often readily be extended to an efficient canonization procedure (c.f., [GN21]). In the case of groups, it is not clear whether WL easily identifies colored variants, which were studied in [BS22, LGR16]. To this end, we ask the following. A colored group is a group $G$ together with a coloring $\chi\colon G\to\mathcal{K}$ for some set of colors $\mathcal{K}$. Two colored groups are isomorphic if there is an isomorphism between them that preserves the colors of all elements. WL can naturally be started with such an initial coloring, which is then refined according to the initial WL coloring as defined above.
@@
For the classes of groups we have studied, when we have been able to give an $O(1)$ bound on their WL-dimension, we also get an $O(\log n)$ bound on the number of rounds needed. The dimension bound alone puts the problem into $\mathsf{P}$, while the bound on rounds puts it into $\mathsf{TC}^{1}$. A priori, these two should be distinct. For example, in the case of graphs, Kiefer & McKay [KM20] have shown that there are graphs for which color refinement takes $n-1$ rounds to stabilize. There has been also considerable work on lower bounds against the iteration number for $k$-WL for graphs, when $k>2$ [Für01, BN23, GLN25].
@@
We also wish to highlight a question that essentially goes back to [CTW13], who showed that GpI cannot be hard under $\mathsf{AC}^{0}$ reductions for any class containing Parity. In Theorem 7.6, we showed that count-free WL requires dimension $\Omega(\log(n))$ (and hence, $\Theta(\log n)$) to identify even Abelian groups. This shows that this particular, natural method does not put GpI into $\mathsf{FO}(\operatorname{poly}\log\log n)$, though it does not actually prove $\textsc{GpI}\notin\mathsf{FO}(\operatorname{poly}\log\log n)$, since we cannot rule out clever bit manipulations of the Cayley (multiplication) tables. While we think the latter lower bound would be of significant interest, we think even the following question is interesting:

### v7

@@ 7.3 Count-free WL and Abelian groups
Case 3: $\langle g^{\prime}_{1}\rangle\cap A_{1}=1$. In this case, $g^{\prime}_{1}$ is part of a basis for $\langle A_{1},g^{\prime}_{1}\rangle$, thus the type of the latter is either $(a+1,b,c)$ if the order of $g^{\prime}_{1}$ is 2, or $(a,b+1,c)$ if the order of $g^{\prime}_{1}$ is 4. In the former case, we have $a+1\leq n-2$ since $G$ only contains $(\mathbb{Z}/2\mathbb{Z})^{n-2}$ as a direct factor, but not any larger power of $(\mathbb{Z}/2\mathbb{Z})$, that is, $G_{n}$ does not contain subgroups of type (n1,,)(n-1,*,*). (And in any case we must have $a+1<n$ since there are only $n$ pebbles.) Thus $a<n$, and by Lemma 7.5 there is an element $h^{\prime}_{1}\in H$ of order 2 that is not a square in $H$ and is not in $B_{1}$, and Duplicator may respond with $h^{\prime}_{1}$. As before, since $\langle h^{\prime}_{1}\rangle\cap B_{1}=1$, there is a unique extension of $\varphi$ to $\langle A_{1},g^{\prime}_{1}\rangle$ that sends $g^{\prime}_{1}\mapsto h^{\prime}_{1}$ which is also good, and thus the new pebbling is good.
@@
Proposition 7.10 (Order finding in WL-I).
@@
Thm. 7.12 is an example of the strategy of using count-free WL, followed by a limited amount of counting afterwards. (We contrast this with the parallel implementation of the classical (counting) WL algorithm, which—for fixed kk—uses a polynomial number of Majority gates at each iteration [GV06].) After the fact, we realized this same bound could be achieved by existing techniques; we include both proofs to highlight an example of how WL was used in the discovery process.
@@
Suppose now that $H$ is Abelian. We run the count-free $(O(1),O(\log\log n))$-WL using the parallel WL implementation due to Grohe & Verbitsky. As $G$ and $H$ are non-isomorphic Abelian groups, they have different order multisets. In particular, there exists an order class of greater multiplicity in $G$ than in $H$. By Prop. 7.10, two elements with different orders receive different colors. We use a $\beta_{1}\textsf{MAC}^{0}$ circuit to distinguish $G$ from $H$. Using $O(\log n)$ non-deterministic bits, we guess the color class $C$ where the multiplicity differs. At each iteration, the parallel WL implementation due to Grohe & Verbitsky records indicators as to whether two $k$-tuples receive the same color. As we have already run the count-free WL algorithm, we may in $\textsf{AC}^{0}$ decide whether two $k$-tuples have the same color. For each $k$-tuple of $G^{k}$ having color class $C$, we feed a $1$ to the Majority gate. For each $k$-tuple of $H^{k}$ having color class $C$, we feed a 0 to the Majority gate. The Majority gate outputs a $1$ if and only if there are strictly more $1$’s than 0’s. The result now follows. ∎
@@ 8 Conclusion
It is often the case that if an uncolored class of graphs is identified by WL, then so is the corresponding class of colored graphs. So if constant-dimensional WL identifies a class of graphs, it can often readily be extended to an efficient canonization procedure (cf., [GN23]). In the case of groups, it is not clear whether WL easily identifies colored variants, which were studied in [BS22, LGR16]. To this end, we ask the following. A colored group is a group $G$ together with a coloring $\chi\colon G\to\mathcal{K}$ for some set of colors $\mathcal{K}$. Two colored groups are isomorphic if there is an isomorphism between them that preserves the colors of all elements. WL can naturally be started with such an initial coloring, which is then refined according to the initial WL coloring as defined above.
@@
For the classes of groups we have studied, when we have been able to give an $O(1)$ bound on their WL-dimension, we also get an $O(\log n)$ bound on the number of rounds needed. The dimension bound alone puts the problem into $\mathsf{P}$, while the bound on rounds puts it into $\mathsf{TC}^{1}$. A priori, these two should be distinct. For example, in the case of graphs, Kiefer & McKay [KM20] have shown that there are graphs for which color refinement takes n1n-1 rounds to stabilize. There has been also considerable work on lower bounds against the iteration number for $k$-WL for graphs, when $k>2$ [Für01, BN23, GLN25].
@@
We also wish to highlight a question that essentially goes back to [CTW13], who showed that GpI cannot be hard under $\mathsf{AC}^{0}$ reductions for any class containing Parity. In Theorem 7.6, we showed that count-free WL requires dimension $\Omega(\log n)$ (and hence, $\Theta(\log n)$) to identify even Abelian groups. This shows that this particular, natural method does not put GpI into $\mathsf{FO}(\operatorname{poly}\log\log n)$, though it does not actually prove $\textsc{GpI}\notin\mathsf{FO}(\operatorname{poly}\log\log n)$, since we cannot rule out clever bit manipulations of the Cayley (multiplication) tables. While we think the latter lower bound would be of significant interest, we think even the following question is interesting: